At 25°C, the following heats of reaction are known: AH (kJ/mol 167.4 2CIF + 02 →Cl20 + F20 2ClF3 + 202 →Cl20+3F20 341.4 2F2 + 02 → 2F20 At the same temperature, calculate ΔH for the reaction: ClF + F2 → CIF3 -43.4 A. -217.5 kJ/mol B.-130.2 kJ/mol C. +217.5 kJ/mol ○ D.-108.7 kJ/mol E. none of these QUESTION 4 Consider the reaction: When a 12.9-g sample of ethyl alcohol(molar mass 46.07 g/mol) is burned, how much energy is released as heat? A. 3.57 kJ OB, 3.84 × 102 kJ -, c. 1.77 × 104 KJ D. 0.384 kJ ○ E. 0.280 kJ QUESTION 5 For a particular process q = 20 kJ and w = 15 kJ, which of the following statements is true? A. Heat flows from the system to the surroundings. B. The system does work on the surroundings. 0C. AE = 35 kJ D. All of the above are true. E. None of the above are true.
media%2Fabc%2Fabce2817-b54d-431d-a0b5-81
Answer #1

Q3

Apply Hes Law

if we need ClF + F2 = ClF3

from the data, invert RXN2 since we need ClF3, divide by 2

divide RXN1 by 2, since we need ClF with 1 coeff.

divide RXN 3 by 2, since we need F2 with 1 stoichiometric coeff

HTotal = -341.4/2 + 1/2*167.4 +1/2*-43.4

Htotal = -108.7 kJ/mol

Q4

mol of alcohol = mass/MW = 12.9/46.07 = 0.280

1 mol o = 1.37*10^3

0.280 mol --> X

x = 0.280*1.37*10^3 = 383.6 kJ = 3.84*10^2 kJ

Q5

Q = +20 kJ --> heat release

W = 15 kJ --> work on system

Q-W = dE

a)

is false, Heat flow TO the system

b)

this is true, system is being expanded

c)

False,

Q-W = dE

20-15 = dE = 5

d and e)

false

choose B